Can someone confirm the solution for me. Using A=P(1+rt)
Find the term of a loan of $1,800 at 9.3% if the simple interest is $2,845.80.
I'm getting t= 6.247 or 6.25 years
Thank you!

Answer by rfer(16322)

If the total interest is $2845.80 total and the interest per year is 1800x0.093=167.40
2845.80/167.40=17 yrs.
If the total interest is 2845.80-1800=1045.80
then 1045.80/167.40=6.25 yrs
Your answer does not fit your question.